can someone help please
Five cards are drawn at random from an ordinary deck of 52 cards. In how many ways is it possible to draw two red cards and three black cards?
a. 845,000 ways
b. 422,500 ways
c. 1,690,000 ways
d. 1,267,500 ways

Answer by stanbon(75887)

Five cards are drawn at random from an ordinary deck of 52 cards. In how many ways is it possible to draw two red cards and three black cards?
---------------
Draw 2 red from 26 red: 26C2 = 13*25 = 325
Draw 3 black from 26 black: 26C3 = 2600
--------------
# of ways to draw 2 red and 3 black = 325*2600 = 845,000
==================================
Cheers,
Stan H.
a. 845,000 ways
b. 422,500 ways
c. 1,690,000 ways
d. 1,267,500 ways